# appReferences Trajectory-aware Shifted State Space Models for Online Video Super-Resolution

### v1

@@ Introduction
Selective Scanning along Temporal Dimension
@@
Loss Function
We adopt Charbonnier loss (Lai et al. 2018) as the spatial loss function to supervise the SR frame generation:
@@
To supervise the trajectory generation for ensuring the accuracy of token selection, we first employ the formulation of trajectories of LR video in Equation (3) to generate trajectories of HR video:
@@
where is the downsampling operation with scale factor that subsamples every coordinate to LR size.
@@
Category |
Methods |
R-T. |
BI degradation | BD degradation | ||||||||||||
Bidirectional |
BasicVSR (Chan et al. 2021) | P+F | ✗ | 63 | 15.9 | 397 | 6.3 | 31.42/0.8909 | 27.24/0.8251 | 37.53/0.9498 | 27.96/0.8553 | ||||||||||
| IconVSR (Chan et al. 2021) | P+F | ✗ | 70 | 14.3 | 452 | 8.7 | 31.67/0.8948 | 27.39/0.8279 | 37.84/0.9524 | 28.04/0.8570 | |||||||||||
| BasicVSR++ (Chan et al. 2022) | P+F | ✗ | 77 | 13.0 | 418 | 7.3 | 32.39/0.9069 | 27.79/0.8400 | 38.21/0.9550 | 29.04/0.8753 | |||||||||||
| SSL-bi (Chan et al. 2021) | P+F | ✗ | 24 | 41.7 | 92 | 1.0 | 31.06/0.8933 | 27.15/0.8208 | 37.06/0.9458 | 27.56/0.8431 | |||||||||||
| Bicubic | N | ✓ | - | - | - | - | 26.14/0.7292 | 23.78/0.6347 | 31.30/0.8687 | 21.80/0.5246 | |||||||||||
Online VSR |
RRN (Chan et al. 2021) | P | ✓ | 34 | 29.4 | 193 | 3.4 | 28.82/0.8234 | 25.85/0.7660 | 36.69/0.9432 | 27.69/0.8488 | ||||||||||
| BasicVSR++* | P | ✓ | 40 | 25.0 | 146 | 3.0 | 30.44/0.8686 | 27.06/0.8173 | 37.11/0.9464 | 27.49/0.8426 | |||||||||||
| DAP-128 (Fuoli et al. 2023) | P | ✓ | 38 | 26.3 | 165 | - | 30.59/0.8703 | - | 37.29/0.9476 | - | |||||||||||
| FDAN (Yang et al. 2023) | P | ✓ | 34 | 29.4 | 146 | 3.9 | 30.71/0.8723 | 27.14/0.8206† | 37.36/0.9483† | 27.76/0.8471 | |||||||||||
| KSNet (Jin et al. 2023) | P | ✓ | 31 | 32.3 | 145 | 3.0 | 30.69/0.8724 | 27.14/0.8208 | 37.34/0.9490 | 27.63/0.8444† | |||||||||||
| TMP (Zhang et al. 2024b) | P | ✓ | 25 | 40.1 | 176 | 3.1 | 30.67/0.8710 | 27.10/0.8167 | 37.33/0.9481 | 27.61/0.8428 | |||||||||||
| TS-Mamba (ours) | P | ✓ | 29 | 33.5 | 112 | 3.0 | 30.73/0.8727 | 27.17/0.8209 | 37.36/0.9482 | 27.70/0.8473 | |||||||||||
Experiments
Experimental Settings
@@
Following the previous online VSR research (Jin et al. 2023; Zhang et al. 2024b), we use REDS (Nah et al. 2019), and Vimeo-90K (Xue et al. 2019) as training datasets. REDS4 is used for evaluating the models trained on the REDS dataset, while Vimeo-90K-T and Vid4 (Liu and Sun 2013) are utilized for benchmarking the models trained on the Vimeo-90K dataset. Two degradations, BI (bicubic) and blur degradation (BD), are used to perform downsampling and the downsampling factor is set to = 4. For BI downsampling, the HR frame is downsampled by a bicubic filter. For BD downsampling, the HR frame is first blurred by a Gaussian filter with standard deviation = 1.6, and then the blurred frame is subsampled for every pixels to generate the LR frame. PSNR and SSIM are adopted as performance evaluation metrics. Runtime (Run.), FPS (frames per second), MACs, and parameters (Params.) are computed on an LR frame of size 180320 to evaluate the model complexity and speed.
@@
In the experiments, the numbers of residual blocks and are set to 2 and 13, respectively. The token size is 44 and the window size is 88. The selected token number is set as 3. Random flips, rotations, and temporal inversion operations are performed for data augmentation. Adam optimizer (Kingma 2014), and Cosine Annealing scheme (Loshchilov and Hutter 2016) are used during network training. The HR patch size is 256256 and the batch size is 8. The total number of iterations is 600K. The proposed method is implemented on the PyTorch platform with two NVIDIA GeForce RTX 3090 GPUs. Following (Liu et al. 2022a), a lightweight optical flow network (Kong, Shen, and Yang 2021) is adopted to update trajectories. The temporal window size is set as 15 based on (Zhang et al. 2024b) when training on REDS (Nah et al. 2019). For the Vimeo-90K (Xue et al. 2019) dataset, the original sequence is temporally flipped to obtain a 14-frame sequence.
@@
We compare our approach with five SOTA online VSR methods, including RRN (Isobe et al. 2020), DAP-128 (Fuoli et al. 2023), FDAN (Yang et al. 2023), KSNet (Jin et al. 2023), and TMP (Zhang et al. 2024b), and four bidirectional propagation VSR methods, BasicVSR (Chan et al. 2021), IconVSR (Chan et al. 2021), BasicVSR++ (Chan et al. 2022), and SSL (Xia et al. 2023)). Additionally, we implemented another method, i.e.,“BasicVSR++*”, by removing the backward propagation branch of BasicVSR++ and reducing its model size for online VSR application. We use “P”,“F” and “N” to represent those with the previous support frames, future support frames and no support frames.
Overall Performance
@@
Following (Fuoli et al. 2023; Zhang et al. 2024b), VSR methods that can process 720p (1280720) videos in at least 24 in terms of FPS are recognized as real-time (R-T.) methods (Fuoli et al. 2023), and we have labelled all the tested methods in Table 1 according to their runtime. It is noted that our TS-Mamba model achieves the second fastest inference speed among all online VSR methods. TMP is the one with the fastest runtime as it was implemented with the CUDA accelerator (high MACs but low runtime) while TS-Mamba is not. Moreover, TS-Mamba also offers a significant reduction in terms of MACs (about 36.3%) and a marginal reduction in parameter numbers compared to TMP, as shown in Figure 1.
Ablation Study
To further verify the effectiveness of our contributions, we have conducted ablation studies on the REDS4 dataset.
| Models | PSNR/SSIM | Params. | Run. | MACs |
@@
We first confirmed the contribution of two trajectory-aware designs, i.e., trajectory generation and trajectory-aware loss, by creating the following variants. (v1.1) w/o Trajectory - and Token Selection module were removed from TS-Mamba; (v1.2) w/o - the trajectory-aware loss function was removed when training the TS-Mamba model. We further verified our proposed TSMA module in terms of compensation branches and shift operations, by obtaining (v1.3) w/o IntraWCB and (v1.4) w/o InterWCB - IntraWCB and InterWCB were removed from TSMA module, respectively; (v1.5) w/o IntraWCB + InterWCB - both IntraWCB and InterWCB were removed from the TSMA module. We also tested the adopted shift operations in compensation branches, and implemented (v1.6) w/o - the shift operations were removed in IntraWCB, (v1.7) w/o - the shift operations were removed in InterWCB and (v1.8) w/o (v1.6)+(v1.7) - all the shift operations were removed in TSMA module. As shown in Table 2, the performance of all these variants is evidently lower than that of the full TS-Mamba, which fully confirmed the effectiveness of each key component in our design.
| PSNR/SSIM | Params. | Run. | MACs | |
@@
To confirm the value of the token number in our TS-Mamba, we tested different values with our TS-Mamba model, and presented the results in Table 3. It is noted that as increases, the VSR performance improves, but with higher model complexity. When = 4, it is difficult to obviously improve VSR performance. To trade off between complexity and performance, we set = 3 in this work.
Conclusion
In this paper, we proposed a Trajectory-aware Shifted SSMs (TS-Mamba) for online VSR, leveraging long-term trajectory modeling and low-complexity Mamba to achieve efficient spatio-temporal information aggregation. In TS-Mamba, trajectories in a video are first constructed to select the most similar tokens from the previous frames. A trajectory-aware shifted Mamba aggregation module is then employed, which consists of shifted SSMs blocks to aggregate the selected tokens. The shifted SSMs blocks are designed based on Hilbert scannings and shift operations to compensate for the scanning losses and strengthen the spatial continuity of Mamba. Moreover, a trajectory-aware loss function is proposed to supervise the trajectory generation, optimizing the accuracy of token selection when training our model. Extensive experiments on three widely used VSR benchmarks have demonstrated the effectiveness and efficiency of our method.

### v2

@@ 3.4 Shifted SSMs Block
3.5 Selective Scanning along Temporal Dimension
@@
3.6 Loss Function
We adopt Charbonnier loss (Lai et al., 2018) as the spatial loss function to supervise the SR frame generation:
@@ 3.6 Loss Function
To supervise the trajectory generation for ensuring the accuracy of token selection, we first employ the formulation of trajectories of LR video in Equation 3 to generate trajectories of HR video:
@@
where is the downsampling with scale factor that subsamples every coordinate to LR size.
@@
4 Experiments
4.1 Experimental Settings
@@
Following previous online VSR research (Jin et al., 2023; Zhang et al., 2024b), we use REDS (Nah et al., 2019), and Vimeo-90K (Xue et al., 2019) as training datasets. REDS4 is used for evaluating the models trained on the REDS dataset, while Vimeo-90K-T and Vid4 (Liu & Sun, 2013) are utilized for benchmarking the models trained on the Vimeo-90K dataset. Two degradations, BI (bicubic) and blur degradation (BD), are used to perform downsampling and the downsampling factor is set to = 4. For BI downsampling, the HR frame is downsampled by a bicubic filter. For BD downsampling, the HR frame is first blurred by a Gaussian filter with standard deviation = 1.6, and then the blurred frame is subsampled for every pixels to generate the LR frame. PSNR and SSIM are adopted as performance evaluation metrics. Runtime (Run.), FPS (frames per second), MACs, and parameters (Params.) are computed on an LR frame of size 180320 to evaluate model complexity and speed.
@@
In the experiments, the numbers of residual blocks and are set to 2 and 13, respectively. The token size is 44 and the window size is 88. The selected token number is set as 3. Random flips, rotations, and temporal inversion operations are performed for data augmentation. Adam optimizer (Kingma, 2014), and Cosine Annealing scheme (Loshchilov & Hutter, 2016) are used during network training. The HR patch size is 256256 and the batch size is 8. The total number of iterations is 600K. The proposed method is implemented on the PyTorch platform with two NVIDIA GeForce RTX 3090 GPUs. Following (Liu et al., 2022a), a lightweight optical flow network (Kong et al., 2021) is adopted to update trajectories. The temporal window size is set as 15 based on (Zhang et al., 2024b) when training on REDS (Nah et al., 2019). For the Vimeo-90K (Xue et al., 2019) dataset, the original sequence is temporally flipped to obtain a 14-frame sequence.
@@
We compare our approach with five SOTA online VSR methods, including RRN (Isobe et al., 2020), DAP-128 (Fuoli et al., 2023), FDAN (Yang et al., 2023), KSNet (Jin et al., 2023), and TMP (Zhang et al., 2024b), and eight bidirectional propagation VSR methods, BasicVSR (Chan et al., 2021), IconVSR (Chan et al., 2021), BasicVSR++ (Chan et al., 2022a), SSL (Xia et al., 2023), DFVSR (Dong et al., 2023), MIA-VSR (Zhou et al., 2024b), IART (Xu et al., 2024) and VSRM (Tran et al., 2025)). Additionally, we implemented another methods, i.e.,“BasicVSR++*” and “VSRM*”, by removing the backward propagation branch of VSR models, i.e., BasicVSR++ and VSRM, and reducing its model size for online VSR application. We use “P”,“F” and “N” to represent those with the previous support (supp.) frames, future support frames and no support frames.
| Category | Methods | Supp. Frame | R-T. | Run. (ms) | FPS (1/s) | MACs (G) | Params. (M) | BI degradation | BD degradation | ||
| REDS4(RGB) (PSNR/SSIM) | Vid4(Y) (PSNR/SSIM) | Vimeo-90K-T(Y) (PSNR/SSIM) | Vid4(Y) (PSNR/SSIM) | ||||||||
| Bidirectional | BasicVSR | P+F | ✗ | 63 | 15.9 | 397 | 6.3 | 31.42/0.8909 | 27.24/0.8251 | 37.53/0.9498 | 27.96/0.8553 |
| IconVSR | P+F | ✗ | 70 | 14.3 | 452 | 8.7 | 31.67/0.8948 | 27.39/0.8279 | 37.84/0.9524 | 28.04/0.8570 | |
| BasicVSR++ | P+F | ✗ | 77 | 13.0 | 418 | 7.3 | 32.39/0.9069 | 27.79/0.8400 | 38.21/0.9550 | 29.04/0.8753 | |
| SSL-bi | P+F | ✗ | 24 | 41.7 | 92 | 1.0 | 31.06/0.8933 | 27.15/0.8208 | 37.06/0.9458 | 27.56/0.8431 | |
| DFVSR | P+F | ✗ | - | - | - | 7.1 | 32.76/0.9081 | 27.92/0.8427 | 38.51/0.9571 | 29.56/0.8983 | |
| MIA-VSR | P+F | ✗ | 318 | 3.1 | 3220 | 16.5 | 32.78/0.9220 | 28.20/0.8507 | - | - | |
| IART | P+F | ✗ | 180 | 5.6 | 5020 | 13.4 | 32.90/0.9138 | 28.26/0.8517 | 38.62/0.9579 | 29.68/0.8884 | |
| VSRM | P+F | ✗ | 223 | 4.5 | 2174 | 17.1 | 33.11/0.9162 | 28.44/0.8552 | - | - | |
| Online | Bicubic | N | ✓ | - | - | - | - | 26.14/0.7292 | 23.78/0.6347 | 31.30/0.8687 | 21.80/0.5246 |
| RRN | P | ✓ | 34 | 29.4 | 193 | 3.4 | 28.82/0.8234 | 25.85/0.7660 | 36.69/0.9432 | 27.69/0.8488 | |
| BasicVSR++* | P | ✓ | 40 | 25.0 | 146 | 3.0 | 30.44/0.8686 | 27.06/0.8173 | 37.11/0.9464 | 27.49/0.8426 | |
| DAP-128 | P | ✓ | 38 | 26.3 | 165 | - | 30.59/0.8703 | - | 37.29/0.9476 | - | |
| FDAN | P | ✓ | 34 | 29.4 | 146 | 3.9 | 30.71/0.8723 | 27.14/0.8206† | 37.36/0.9483† | 27.76/0.8471 | |
| KSNet | P | ✓ | 31 | 32.3 | 145 | 3.0 | 30.69/0.8724 | 27.14/0.8208 | 37.34/0.9490 | 27.63/0.8444† | |
| TMP | P | ✓ | 25 | 40.1 | 176 | 3.1 | 30.67/0.8710 | 27.10/0.8167 | 37.33/0.9481 | 27.61/0.8428 | |
| VSRM* | P | ✓ | 31 | 32.7 | 136 | 3.1 | 30.64/0.8701 | 27.10/0.8163 | 37.28/0.9477 | 27.57/0.8423 | |
| TS-Mamba | P | ✓ | 29 | 33.5 | 112 | 3.0 | 30.73/0.8727 | 27.17/0.8209 | 37.36/0.9482 | 27.70/0.8473 | |
4.2 Overall Performance
@@
Following (Fuoli et al., 2023; Zhang et al., 2024b), VSR methods that can process 720p (1280720) videos in at least 24 in terms of FPS are recognized as real-time (R-T.) methods (Fuoli et al., 2023), and we have labelled all the tested methods in Table 1 according to their runtime. It is noted that our TS-Mamba model achieves the second fastest inference speed among all online VSR methods. TMP is the one with the fastest runtime as it was implemented with the CUDA accelerator (high MACs but low runtime) while TS-Mamba is not. Moreover, TS-Mamba also offers a significant reduction in terms of MACs (about 36.3%) and a marginal reduction in parameter numbers compared to TMP, as shown in Figure 1.
4.3 Ablation Study
To further verify the effectiveness of our contributions, we have conducted ablation studies on the REDS4 dataset with BI degradation.
| Models | PSNR(dB) / SSIM | Params.(M) | Run.(ms) | MACs(G) |
@@ 4.3 Ablation Study
| GT | w/o Trajectory | w/o | w/o IntraWCB | w/o InterWCB | |
| REDS4_000_045 (BI) | w/o IntraWCB + InterWCB | w/o | w/o | w/o (v1.6)+(v1.7) | TS-Mamba (ours) |
We first confirmed the contribution of two trajectory-aware designs, i.e., trajectory generation and trajectory-aware loss, by creating the following variants. (v1.1) w/o Trajectory - and Token Selection module were removed from TS-Mamba; (v1.2) w/o - the trajectory-aware loss function was removed when training the TS-Mamba model. We further verified our proposed TSMA module in terms of compensation branches and shift operations, by obtaining (v1.3) w/o IntraWCB and (v1.4) w/o InterWCB - IntraWCB and InterWCB were removed from the TSMA module, respectively; (v1.5) w/o IntraWCB + InterWCB - both IntraWCB and InterWCB were removed from the TSMA module. We also tested the adopted shift operations in compensation branches, and implemented (v1.6) w/o - the shift operations were removed in IntraWCB, (v1.7) w/o - the shift operations were removed in InterWCB and (v1.8) w/o (v1.6)+(v1.7) - all the shift operations were removed in TSMA module. As shown in Table 2, the performance of all these variants is evidently lower than that of the full TS-Mamba, which fully confirms the effectiveness of each key component in our design. Moreover, we further provide the visual results of these variants on REDS4 dataset in Figure 5. It is found that the visual results demonstrates the contributions of our designs, particularly in realistic textures and fine details of the car.
| PSNR(dB)/SSIM | Params. (M) | Run. (ms) | MACs (G) | |
@@
REDS4_015_049
GT
FDAN
TMP
VSRM*
TS-Mamba
To confirm the value of the token number in our TS-Mamba, we tested different values with our TS-Mamba, and presented the results in Table 3. It is noted that as increases, the VSR performance improves, but with higher model complexity. When = 4, it is difficult to obviously improve VSR performance. To trade off between complexity and performance, we set = 3 in this work.
5 Limitations
We investigate our results and find out the failure cases. A failure case when highly dynamic rotation occurs is visualized in Figure 6. The generated trajectories and compensated manner of our TS-Mamba are inaccurate enough when dynamic rotation occurs in car tire and rotation information cannot be reconstructed, thus limiting the performance of our method. Due to the high difficulty of modeling rotation, other online VSR methods also fail to obtain complete rotation information.
6 Conclusion
In this paper, we proposed a Trajectory-aware Shifted SSMs (TS-Mamba) for online VSR, leveraging long-term trajectory modeling and low-complexity Mamba to achieve efficient spatio-temporal information aggregation. In TS-Mamba, trajectories in a video are first constructed to select the most similar tokens from the previous frames. A trajectory-aware shifted Mamba aggregation module is then employed, which consists of shifted SSMs blocks to aggregate the selected tokens. The shifted SSMs blocks are designed based on Hilbert scannings and shift operations to compensate for the scanning losses and strengthen the spatial continuity of Mamba. Extensive experiments on three widely used VSR benchmarks have demonstrated the effectiveness and efficiency of our method.